OpenStudy (anonymous):

arnold has taken 4 tests. on the 1st test he scored 83%, on the 2nd he got 85%, and on the 3rd he got 100%. the 4th test is counted double any of the other tests. what score did he get on the 4th test if his final overall average was 90%

OpenStudy (anonymous):

90%= 0.9 90=(83+85+100+x)/4 90=(268+X)/4 (90*4)=268+x 360 -268=x x = 92% \[4th test \ge than 92 percent\]

OpenStudy (anonymous):

92%